# Stochastic Cahn–Hilliard Equations from One-Dimensional Ising–Kac–Kawasaki Dynamics

### 31 Mar 2026 v1

@@ 5.1 Convergence of the discrete martingale
Since $\varepsilon_{\gamma}^{2}/\delta_{\gamma}\to 0$, the jumps vanish. Martingale central limit theorem yields
$M^{\gamma}(\phi)\Longrightarrow M(\phi)$
in $D([0,T],\mathbb{R})$, where $M(\phi)$ is a continuous centered Gaussian martingale with quadratic variation
@@ 5.2 Convergence of the stochastic convolution in $\mathcal{C}^{\alpha}$
Under the Assumption $({\bf A})$ , for $k,\ell\in\mathcal{B}_{\gamma}$,
@@
$(\mathcal{K}_{\gamma}^{\varepsilon}e_{k})(\varepsilon_{\gamma}i)=\varepsilon_{\gamma}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)e^{2\pi ik\varepsilon_{\gamma}j}=\varepsilon_{\gamma}\left(\sum_{m\in\Lambda_{N}}\kappa_{\gamma}(m)e^{-2\pi ik\varepsilon_{\gamma}m}\right)e^{2\pi ik\varepsilon_{\gamma}i}.$
@@
$\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}e_{k})(\varepsilon_{\gamma}i)=\varepsilon_{\gamma}\left(\sum_{m\in\Lambda_{N}}\kappa_{\gamma}(m)e^{-2\pi ik\varepsilon_{\gamma}m}\right)\frac{e_{k}(\varepsilon_{\gamma}(i+1))-e_{k}(\varepsilon_{\gamma}i)}{\varepsilon_{\gamma}}.$
@@
$\left\langle\widehat{M}_{\gamma}(k),\overline{\widehat{M}_{\gamma}(\ell)}\right\rangle_{t}=\widetilde{\sigma}_{\gamma}^{2}\,t\,\vartheta_{\gamma}(-k)D_{\gamma}(-k)\,\overline{\vartheta_{\gamma}(-\ell)D_{\gamma}(-\ell)}\,\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}e_{-k}(\varepsilon_{\gamma}i)e_{\ell}(\varepsilon_{\gamma}i)+o_{\gamma}(1).$
@@
$\sup_{\gamma>0}\mathbb{E}\Bigl[\sup_{t\leq T}\|\widetilde{Z}_{\gamma}(t)\|_{B^{\beta}_{p,p}}^{p}\Bigr]=$ $\sup_{\gamma>0}\mathbb{E}\Bigl[\sup_{t\leq T}2^{\beta pq}\|\Delta_{j}\widetilde{Z}_{\gamma}(t)\|_{L^{p}}^{p}\Bigr]$
$\leq$ $\sum_{q\geq-1}2^{\beta pq}2^{-qp/2}<\infty,$ (36)
@@
Applying Burkholder–Davis–Gundy inequality together with Lemma 13 and semigroup group estimates 4, we obtain
@@
For every $\alpha<\frac{1}{2}$, the discrete stochastic convolution $\widetilde{Z}_{\gamma}\Longrightarrow Z$ in $C([0,T],\mathcal{C}^{\alpha}(\mathbb{T})).$
@@
Lemma 14 yields that, passing to a further subsequence if necessary, we have
$\widetilde{Z}_{\gamma_{n}}\Longrightarrow\widetilde{Z}\qquad\text{in }C([0,T],B^{\beta}_{p,p}(\mathbb{T})).$
To identify the limit, it is enough to show that every convergent subsequence has limit equal to $Z$. We now prove that $\widetilde{Z}=Z$ in law.
@@
It remains to control the high-frequency tail. By Step 1,
@@ 6 Energy estimates for $Y_{\gamma}$
In this section, we show the uniform discrete $H^{-1}$-energy estimate for $Y_{\gamma}$.
@@
Assume that $Y_{\gamma}$ solves
$\partial_{t}Y_{\gamma}=-\nu_{\gamma}\Delta_{\varepsilon_{\gamma}}^{2}Y_{\gamma}-A_{\gamma}\Delta_{\varepsilon_{\gamma}}(Y_{\gamma}+Z_{\gamma})+\frac{\beta\lambda_{\gamma}}{6}\Delta_{\varepsilon_{\gamma}}\bigl((Y_{\gamma}+Z_{\gamma})^{3}\bigr)+R_{\gamma},\quad Y_{\gamma}(0)=X_{\gamma}(0),$
on the lattice subspace $\Lambda_{\varepsilon_{\gamma}}$. Then there exist constants $C_{T}>0$, independent of $\gamma$, such that for every $T>0$,
$\sup_{t\in[0,T]}\|Y_{\gamma}(t)\|_{-1,\gamma}^{2}+$ $\int_{0}^{T}\|Y_{\gamma}(s)\|_{1,\gamma}^{2}\,ds+\int_{0}^{T}\|Y_{\gamma}(s)\|_{L^{4}_{\gamma}}^{4}\,ds$
$\leq C_{T}\left(\|X_{\gamma}(0)\|_{-1,\gamma}^{2}+1+\int_{0}^{T}\|Z_{\gamma}(s)\|_{L^{4}_{\gamma}}^{4}\,ds+\int_{0}^{T}\|R_{\gamma}(s)\|_{-1,\gamma}^{2}\,ds\right).$ (37)
@@
Testing the equation against $(-\Delta_{\varepsilon_{\gamma}})^{-1}Y_{\gamma}$, we obtain
$\frac{1}{2}\nu_{\gamma}\frac{d}{dt}\|Y_{\gamma}\|_{-1,\gamma}^{2}+\|Y_{\gamma}\|_{1,\gamma}^{2}=-A_{\gamma}\langle Y_{\gamma}+Z_{\gamma},Y_{\gamma}\rangle_{\gamma}-\frac{\beta\lambda_{\gamma}}{6}\langle(Y_{\gamma}+Z_{\gamma})^{3},Y_{\gamma}\rangle_{\gamma}+\langle R_{\gamma},(-\Delta_{\varepsilon_{\gamma}})^{-1}Y_{\gamma}\rangle_{\gamma}.$
For the linear term, by the discrete Poincaré inequality and Young’s inequality,
$|\langle Y_{\gamma}+Z_{\gamma},Y_{\gamma}\rangle_{\gamma}|\leq\eta\|Y_{\gamma}\|_{1,\gamma}^{2}+C_{\eta}\|Z_{\gamma}\|_{L^{2}_{\gamma}}^{2}.$
@@
$\frac{\beta\lambda_{\gamma}}{6}\langle(Y_{\gamma}+Z_{\gamma})^{3},Y_{\gamma}\rangle_{\gamma}\leq-c_{1}\|Y_{\gamma}\|_{L^{4}_{\gamma}}^{4}+C\|Z_{\gamma}\|_{L^{4}_{\gamma}}^{4}.$
Finally,
$|\langle R_{\gamma},(-\Delta_{\varepsilon_{\gamma}})^{-1}Y_{\gamma}\rangle_{\gamma}|\leq\frac{1}{2}\|Y_{\gamma}\|_{-1,\gamma}^{2}+\frac{1}{2}\|R_{\gamma}\|_{-1,\gamma}^{2}.$
Collecting the three estimates and choosing $\eta>0$ sufficiently small yields
$\frac{1}{2}\frac{d}{dt}\|Y_{\gamma}\|_{-1,\gamma}^{2}+c\|Y_{\gamma}\|_{1,\gamma}^{2}+c\|Y_{\gamma}\|_{L^{4}_{\gamma}}^{4}\leq C\Bigl(1+\|Z_{\gamma}\|_{L^{4}_{\gamma}}^{4}+\|Y_{\gamma}\|_{-1,\gamma}^{2}+\|R_{\gamma}\|_{-1,\gamma}^{2}\Bigr).$
@@ 7 Convergence to the stochastic Cahn–Hilliard equation
In this section, we proof the convergence in the main result.
uniformly in $\gamma$.
Note that $\widetilde{Y}_{\gamma}$ satisfies, in weak form,
$\partial_{t}\widetilde{Y}_{\gamma}=-\nu_{\gamma}\Delta^{2}\widetilde{Y}_{\gamma}-A_{\gamma}\Delta(\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})+\frac{\beta\lambda_{\gamma}}{6}\Delta\bigl((\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})^{3}\bigr)+R_{\gamma},$
Now we show the convergence of $X_{\gamma}$ to the stochastic Cahn–Hilliard limit.
@@
Assume that the initial data converge: $\widetilde{X}_{\gamma}(0)\rightarrow X_{0},\text{in }H^{-1}(\mathbb{T});$ Then as $\gamma\to 0$, $\widetilde{X}_{\gamma}$ converges inlaw to $X$ in $L^{2}(0,T;L^{2}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T})),$ where $X=Y+Z$ is the unique weak solution of the stochastic Cahn–Hilliard equation.
@@
$\partial_{t}Y=-\nu\Delta^{2}Y+\frac{(1+\beta)\lambda}{2}\Delta(Y+Z)+\frac{\beta\lambda}{6}\Delta((Y+Z)^{3}),\quad Y(0)=X_{0},$
in weak sense. By the uniqueness of the shifted equation, then all subsequences have the same limit, and therefore the full sequence converges: $\widetilde{Y}_{\gamma}\Longrightarrow Y$ in $L^{2}(0,T;H^{1-\kappa}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T})),\forall\kappa\in(0,1).$ Therefore,
@@
$dX=\Bigl(-\nu\Delta^{2}X-\frac{(1+\beta)\lambda}{2}\Delta X+\frac{\beta\lambda}{6}\Delta(X^{3})\Bigr)\,dt+\sigma_{*}\partial_{x}dW(t,x),\qquad X(0)=X_{0}.$

### 05 Apr 2026 v2

@@ 5.1 Convergence of the discrete martingale
Since $\varepsilon_{\gamma}^{2}/\delta_{\gamma}\to 0$, the jumps vanish. Martingale central limit theorem yields $M^{\gamma}(\phi)\rightarrow M(\phi)$ in $D([0,T],\mathbb{R})$, where $M(\phi)$ is a continuous centered Gaussian martingale with quadratic variation
@@ 5.2 Convergence of the stochastic convolution in $\mathcal{C}^{\alpha}$
Under the Assumption $({\bf A})-({\bf C})$ , for $k,\ell\in\mathcal{B}_{\gamma}$,
@@
$(\mathcal{K}_{\gamma}^{\varepsilon}e_{k})(\varepsilon_{\gamma}i)=\varepsilon_{\gamma}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)e^{2\pi ik\varepsilon_{\gamma}j}=\left(\sum_{m\in\Lambda_{N}}\kappa_{\gamma}(m)e^{-2\pi ik\varepsilon_{\gamma}m}\right)e^{2\pi ik\varepsilon_{\gamma}i}.$
@@
$\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}e_{k})(\varepsilon_{\gamma}i)=\vartheta_{\gamma}(k)\,\frac{e_{k}(\varepsilon_{\gamma}(i+1))-e_{k}(\varepsilon_{\gamma}i)}{\varepsilon_{\gamma}}=\vartheta_{\gamma}(k)\,D_{\gamma}(k)\,e_{k}(\varepsilon_{\gamma}i),$
@@
$\left\langle\widehat{M}_{\gamma}(k),\overline{\widehat{M}_{\gamma}(\ell)}\right\rangle_{t}$ $=\widetilde{\sigma}_{\gamma}^{2}\,t\,\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}e_{-k})(\varepsilon_{\gamma}i)\,\overline{\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}e_{-\ell})(\varepsilon_{\gamma}i)}+o_{\gamma}(1)$
$=\widetilde{\sigma}_{\gamma}^{2}\,t\,\vartheta_{\gamma}(-k)D_{\gamma}(-k)\,\overline{\vartheta_{\gamma}(-\ell)D_{\gamma}(-\ell)}\,\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}e_{-k}(\varepsilon_{\gamma}i)e_{\ell}(\varepsilon_{\gamma}i)+o_{\gamma}(1).$
@@
$\sup_{\gamma>0}\mathbb{E}\Bigl[\sup_{t\leq T}\|\widetilde{Z}_{\gamma}(t)\|_{B^{\beta}_{p,p}}^{p}\Bigr]=$ $\sup_{\gamma>0}\mathbb{E}\Bigl[\sup_{t\leq T}2^{\beta pj}\sum_{j\geq-1}\|\Delta_{j}\widetilde{Z}_{\gamma}(t)\|_{L^{p}}^{p}\Bigr]$
$\leq$ $\sum_{j\geq-1}2^{\beta pj}2^{-jp/2}<\infty,$ (44)
@@
Applying Burkholder–Davis–Gundy inequality together with Lemma 14 and semigroup estimates 4, we obtain
@@
For every $\alpha<\frac{1}{2}$, the discrete stochastic convolution $\widetilde{Z}_{\gamma}\rightarrow Z$ in $C([0,T],\mathcal{C}^{\alpha}(\mathbb{T})).$
@@
Lemma 15 yields that, passing to a further subsequence if necessary, we have $\widetilde{Z}_{\gamma_{n}}\rightarrow\widetilde{Z}$ in $C([0,T],B^{\beta}_{p,p}(\mathbb{T}))$. To identify the limit, it is enough to show that every convergent subsequence has limit equal to $Z$. We now prove that $\widetilde{Z}=Z$ in law.
@@
It remains to control the high-frequency tail. By Lemma 15,
@@ 6 Energy estimates for $Y_{\gamma}$
In this section, we show the uniform $H^{-1}$-energy estimate for $Y_{\gamma}$.
uniformly in $\gamma$. Thus $\widetilde{Y}_{\gamma}$ satisfies, in weak form,
$\partial_{t}\widetilde{Y}_{\gamma}=-\nu_{\gamma}\Delta^{2}\widetilde{Y}_{\gamma}-A_{\gamma}\Delta(\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})+\frac{\beta\lambda_{\gamma}}{6}\Delta\bigl((\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})^{3}\bigr)+R_{\gamma},$ (45)
@@
There exist constants $C_{T}>0$, independent of $\gamma$, such that for every $T>0$,
$\sup_{t\in[0,T]}\|\widetilde{Y}_{\gamma}(t)\|_{H^{-1}(\mathbb{T})}^{2}+$ $\int_{0}^{T}\|\widetilde{Y}_{\gamma}(s)\|_{H^{1}(\mathbb{T})}^{2}\,ds+\int_{0}^{T}\|\widetilde{Y}_{\gamma}(s)\|_{L^{4}(\mathbb{T})}^{4}\,ds$
$\leq CT\left(1+\|\widetilde{X}_{\gamma}(0)\|_{H^{-1}(\mathbb{T})}^{2}+\int_{0}^{T}\|\widetilde{Z}_{\gamma}(s)\|_{L^{4}(\mathbb{T})}^{4}\,ds\right).$ (46)
@@
Testing the equation (45) against $(-\Delta)^{-1}\widetilde{Y}_{\gamma}$, we obtain
$\frac{1}{2}\nu_{\gamma}\frac{d}{dt}\|\widetilde{Y}_{\gamma}\|_{H^{-1}(\mathbb{T})}^{2}+\|\widetilde{Y}_{\gamma}\|_{H^{1}(\mathbb{T})}^{2}=-A_{\gamma}\langle\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma},Y_{\gamma}\rangle-\frac{\beta\lambda_{\gamma}}{6}\langle(\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})^{3},\widetilde{Y}_{\gamma}\rangle+\langle R_{\gamma},(-\Delta)^{-1}\widetilde{Y}_{\gamma}\rangle.$
For the linear term, by the Poincaré inequality and Young’s inequality,
$|\langle Y_{\gamma}+Z_{\gamma},Y_{\gamma}\rangle|\leq\eta\|Y_{\gamma}\|_{H^{1}(\mathbb{T})}^{2}+C_{\eta}\|Z_{\gamma}\|_{L^{2}}^{2}.$ (48)
@@
$\frac{\beta\lambda_{\gamma}}{6}\langle(\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})^{3},\widetilde{Y}_{\gamma}\rangle\leq-c_{1}\|\widetilde{Y}_{\gamma}\|_{L^{4}(\mathbb{T})}^{4}+C\|\widetilde{Z}_{\gamma}\|_{L^{4}(\mathbb{T})}^{4}.$ (49)
Finally, since $\frac{1}{2}\|R_{\gamma}\|_{-1,\gamma}^{2}\to 0$ as $\gamma\to 0$,
$|\langle R_{\gamma},(-\Delta)^{-1}\widetilde{Y}_{\gamma}\rangle|\leq\frac{1}{2}\|Y_{\gamma}\|_{H^{-1}}^{2}+1.$ (50)
Collecting the estimates (48)-(50), and choosing $\eta>0$ sufficiently small yields
$\frac{1}{2}\frac{d}{dt}\|Y_{\gamma}\|_{-1,\gamma}^{2}+c\|Y_{\gamma}\|_{1,\gamma}^{2}+c\|Y_{\gamma}\|_{L^{4}_{\gamma}}^{4}\leq C\Bigl(1+\|Z_{\gamma}\|_{L^{4}_{\gamma}}^{4}+\|Y_{\gamma}\|_{-1,\gamma}^{2}\Bigr).$
@@ 7 Convergence to the stochastic Cahn–Hilliard equation
In this section, we proof the convergence of $X_{\gamma}$ to the stochastic Cahn–Hilliard limit. Then we show that the induced canonical Gibbs measure $\mu_{\gamma}:=(X_{\gamma})_{\#}\mu_{N,\gamma,\beta}$ converge to the canonical $\phi^{4}_{1}$ measure on the conserved-mass hyperplane $V_{M}$ in distribution sense.
@@
Assume that the initial data converge: $\widetilde{X}_{\gamma}(0)\rightarrow X_{0},\text{in }H^{-1}(\mathbb{T});$ Then as $\gamma\to 0$, $\widetilde{X}_{\gamma}$ converges inlaw to $X$ in $L^{2}(0,T;L^{2}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T})),$ where $X=Y+Z\in V_{M}$ is the unique weak solution of the stochastic Cahn–Hilliard equation on the conserved-mass hyperplane $V_{M}:=\{\phi\in L^{2}:\langle 1,\phi\rangle=M\}$.
@@
$\partial_{t}Y=-\nu\Delta^{2}Y-\frac{(1+\beta)\lambda}{2}\Delta(Y+Z)+\frac{\beta\lambda}{6}\Delta((Y+Z)^{3}),\quad Y(0)=X_{0},$
in weak sense. By the uniqueness of the shifted equation, then all subsequences have the same limit, and therefore the full sequence converges: $\widetilde{Y}_{\gamma}\rightarrow Y$ in $L^{2}(0,T;H^{1-\kappa}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T})),\forall\kappa\in(0,1).$ Therefore,
@@
$\partial_{t}X=-\nu\Delta^{2}X-\frac{(1+\beta)\lambda}{2}\Delta X+\frac{\beta\lambda}{6}\Delta(X^{3})+\sigma_{*}\nabla\xi,\qquad X(0)=X_{0}.$
@@
Let $\nu_{\gamma}:=(X_{\gamma})_{\#}\mu_{N,\gamma,\beta}$ be the equilibrium measure of the Kac coarse-gained field $\widetilde{X}_{\gamma}$ induced by the microscopic canonical Gibbs measure $\mu_{N,\gamma,\beta}$ on the fixed total-magnetization sector. By construction of $X_{\gamma}$, $\nu_{\gamma}$ is supported on the same conserved-mass hyperplane as $X_{\gamma}$. By above convergence result, we also have the convergence of canonical Gibbs measures.
Proposition 14.
Define the induced equilibrium measure of $X_{\gamma}$ by $\mu_{\gamma}:=(X_{\gamma})_{\#}\mu_{N,\gamma,\beta}.$ Then $\mu_{\gamma}\rightarrow\mu$ in distribution as $\gamma\to 0$, where $\mu$ is the $\phi^{4}_{1}$ measure on conserved-mass hyperplane $V_{M}\}$ defined by (3) with coefficients $A=\frac{(1+\beta)\lambda}{2}$ and $\chi=2\lambda\beta$.
Proof.
Since the Kawasaki dynamics is reversible, and invariant, with respect to the canonical Gibbs measure $\mu_{N,\gamma,\beta}$, if the initial law $\mathrm{Law}\bigl(\sigma^{N}(0)\bigr)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\mu_{N,\gamma,\beta}$, then
$\mathrm{Law}\bigl(\sigma^{N}(t)\bigr)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\mu_{N,\gamma,\beta}\qquad\text{for every }t\geq 0.$
Applying the measurable map $\sigma^{N}\mapsto X_{\gamma}(\sigma)$, we obtain the stationary measure of $X_{\gamma}$, which is given by
$\mathrm{Law}\bigl(X_{\gamma}(t)\bigr)=(X_{\gamma})_{\#}\mathrm{Law}\bigl(\sigma_{\gamma}(t)\bigr)=(X_{\gamma})_{\#}\mu_{N,\gamma,\beta}=\mu_{\gamma}.$
Now we show the tightness of $\{\mu_{\gamma}\}$. By boundedness of $\|\widetilde{Z}\|_{\mathscr{B}^{-\alpha}_{p,q}}$, and the convergence of $\widetilde{Z}_{\gamma}$, for every $\alpha\in(1,1/2)$ and $\gamma>0$, we have
$\frac{1}{T}\int_{0}^{T}\mathbb{E}[\|\widetilde{Z}_{\gamma}(t)\|_{C^{\alpha}}]dt\leq C_{1}$ (51)
for some $C_{1}>0$ uniform in $\gamma$ and $T$. Then using the uniform discrete $H^{-1}$-energy estimate (12) and Soblowv embedding $H^{1}(\mathbb{T})\hookrightarrow C^{\frac{1}{2}}(\mathbb{T})$, we have
$\frac{1}{T}\int_{0}^{T}\mathbb{E}\left[\|\widetilde{Y}_{\gamma}(t)\|_{C^{\frac{1}{2}}}^{2}\right]dt\leq\frac{C}{T}\int_{0}^{T}\mathbb{E}[\|\widetilde{Y}_{\gamma}(t)\|_{H^{1}}^{2}]dt\leq C_{2}+\frac{C_{2}}{T}\mathbb{E}[\|\widetilde{X}_{\gamma}(0)\|_{H^{-1}}^{2}],$ (52)
for some $C_{2}>0$ uniform in $\gamma$ and $T$.
Combining with (51) and (52), the stochastic field $\widetilde{X}_{\gamma}(t)=\widetilde{Y}_{\gamma}(t)+\widetilde{Z}_{\gamma}(t)$ satisfies
$\frac{1}{T}\int_{0}^{T}\mathbb{E}[\|\widetilde{X}_{\gamma}(t)\|_{C^{\alpha}}^{2}]dt\leq C_{3}+\frac{C_{3}}{T}\mathbb{E}[\|\widetilde{X}_{\gamma}(0)\|_{H^{-1}}^{2}],$
for every $\alpha\in(0,\frac{1}{2})$, and some $C_{3}>0$ which is independent with $\gamma$ and $T$. Since $\mathrm{Law}\bigl(X_{\gamma}(t)\bigr)\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\mu_{\gamma}$, we can let $T\to\infty$ to obtain that $\mathbb{E}^{\mu_{\gamma}}[\|X_{\gamma}\|_{C^{\alpha}}^{2}]\leq C_{4}$, which is uniform in $\gamma$. Note that the embedding $C^{\alpha}\hookrightarrow C^{\alpha^{\prime}}$ is compact if $\alpha<\alpha^{\prime}$. Therefore, $\{\mu_{\gamma}\}_{\gamma}$ is tight on $C^{\alpha}$ for every $\alpha\in(0,1/2)$.
Now we fix a subsequence $\gamma_{n}\downarrow 0$ such that $\mu_{\gamma_{n}}\rightarrow\mu_{*}.$ in distribution sense. By Skorokhod’s representation theorem, there are $C^{\alpha}$ value random variables $\{\widetilde{X}_{\gamma_{n}}(0)\}_{n\geq 1}$ and $X(0)$, so that $\mathrm{Law}(\widetilde{X}_{\gamma_{n}}(0))=\mu_{\gamma_{n}}$, $\mathrm{Law}(\widetilde{X}(0))=\mu_{*}$, and $\widetilde{X}_{\gamma_{n}}(0)\to X(0)$ almost surely in $C^{\alpha}$ as $\gamma_{n}\to 0$.
Since each $\widetilde{X}_{\gamma_{n}}$ is stationary and $\widetilde{X}_{\gamma}\to X$ in distribution, for every bounded continuous $F:L^{2}(\mathbb{T})\to\mathbb{R}$ and every $t\in[0,T]$, we pass to the limit as $\gamma_{n}\to 0$ and obtain
$\lim_{\gamma_{n}\to 0}\mathbb{E}\bigl[F(\widetilde{X}_{\gamma_{n}}(t))\bigr]=\lim_{\gamma_{n}\to 0}\int_{C^{\alpha}(\mathbb{T})}F(\phi)\,\mu_{\gamma_{n}}(d\phi)=\int_{C^{\alpha}(\mathbb{T})}F(\phi)\,\mu_{*}(d\phi)=\mathbb{E}\bigl[F(X(t))\bigr]$
for every $t\in[0,T]$. Thus $\mathrm{Law}(X(t))=\mu_{*}$ for every $t$, i.e. the limit process $X$ is stationary. Therefore $\mu_{*}$ is an invariant measure for the limiting stochastic Cahn–Hilliard equation. By the conservation-law, the limit measure $\mu_{*}$ is supported on the conserved-mass hyperplane $V_{M}:=\{\phi\in L^{2}:\langle 1,\phi\rangle=M\}$. Since the limiting conservative stochastic Cahn–Hilliard equation admits a unique $\phi^{4}_{1}$ Gibbs measure $\mu$ on the prescribed mass hyperplane $V_{M}$, every subsequence of $\{\nu_{\gamma}\}$ has the same limit measure $\mu$. Therefore the full sequence $\mu_{\gamma}$ converges weakly to the conserved-mass $\phi^{4}_{1}$ measure. This completes the proof. ∎